A pair of point charges attract each other through Coulumb's interaction with a force of 215 N when they are placed at a distance of 150 m apart. If the first charge increases by a factor of 3; the second charge decreases by a factor of 6, and the distance decreases by a factor of 4. What is the new force?
Answer #1

Solution:

Using the coulomb's law,

=> F = k q1 q2 / r^2

=> new force , F' = k q1' q2' / r'^2

=> F' = (k * 3 * q1 * q2/6 )/(r/4)^2

=> F' = 3 * 16 / 6 * k * q1 * q2 / r^2

=> F' = 8 * k * q1 * q2 / r^2

=> F' = 8 * F

=> F' = 8 * 215

=> F' = 1720 N

please comment for queries.

thanks.